# Quasiparticle interference and spectral function of the UTe2 superconductive surface band

### v1

@@ II.3 Surface Green function calculation
The $T$-matrix method is an exact analytical method to compute the Green’s function for a $\delta$-localized impurity in a non-interacting infinite system [33]. This method can be extended to compute the boundary states of a $d$-dimensional bulk Hamiltonian by considering a ($d$-1)-dimensional localized impurity, which effectively splits the system in half [31, 32] for a very large impurity potential. For example, for a 3D system, the Green’s function computed near a plane-like impurity will converge to the surface Green function in the limit of a large impurity potential.
@@
$G_{\mathrm{BdG}}(E,\mathbf{k})=\Big{(}E+i\eta-\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})\Big{)}^{-1},$ (14)
[…]
$\mathds{V}=V_{0}\begin{pmatrix}\mathds{1}_{8}&0\\ 0&-\mathds{1}_{8}\\ \end{pmatrix}$ (15)
@@
$T(E,\mathbf{k}_{\parallel})=\left(\mathds{1}_{16}-\mathds{V}\int_{BZ_{\perp}}\dfrac{\mathrm{d}\mathbf{k}_{\perp}}{L_{BZ_{\perp}}}G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp},\mathbf{k}_{\parallel})\right)^{-1}\mathds{V}$
[…]
@@
$+G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp 1},\mathbf{k}_{\parallel})T(E,\mathbf{k}_{\parallel})G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp 2},\mathbf{k}_{\parallel})$ (17)
@@
×$\times G(E,\mathbf{k}_{\perp 1},\mathbf{k}_{\perp 2},\mathbf{k}_{\parallel})$
with $d_{\perp}=bc/\sqrt{b^{2}+c^{2}}$ is the distance between the (0-11) impurity plane to its nearest lattice plane. Using Eq. II.3, we can decompose $G_{s}(E,\mathbf{k}_{\parallel})$ into a sum of two terms
$G_{s}(E,\mathbf{k}_{\parallel})=G_{b}(E,\mathbf{k}_{\parallel},0)+G_{i}(E,\mathbf{k}_{\parallel})$ (19)
@@
$G_{b}(E,\mathbf{k}_{\parallel},z_{\perp})$ $=$ $\int_{BZ_{\perp}}\dfrac{\mathrm{d}\mathbf{k}_{\perp}}{L_{BZ_{\perp}}}e^{iz_{\perp}\mathbf{k}_{\perp}\cdot\mathbf{e}_{\perp}}G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp},\mathbf{k}_{\parallel})$
$G_{i}(E,\mathbf{k}_{\parallel})$ $=$ $G_{b}(E,\mathbf{k}_{\parallel},d_{\perp})T(E,\mathbf{k}_{\parallel})G_{b}(E,\mathbf{k}_{\parallel},-d_{\perp})$
We note that the bulk contribution corresponds solely to a simple projection of the 3D-bulk physics on the (0-11) surface, without taking into account the existence of a surface and the semi-infinite nature of the system. The surface physics is hidden in the impurity term $G_{i}(E,\mathbf{k}_{\parallel})$ which when the impurity potential is taken to infinity will describe the formation of the surface states.
@@
$A_{s}(E,\mathbf{k}_{\parallel})=-\frac{1}{\pi}\mathrm{Im}\{\mathrm{Tr}_{\mathrm{el}}[G_{s}(E,\mathbf{k}_{\parallel})]\}$ (22)
where the trace runs over the electron bands. We can also here distinguish two components, and in what follows we will sometimes calculate separately the “bulk” spectral function
$A_{b}(E,\mathbf{k}_{\parallel})=-\frac{1}{\pi}\mathrm{Im}\{\mathrm{Tr}_{\mathrm{el}}[G_{b}(E,\mathbf{k}_{\parallel},0)]\},$ (23)
which is just the component of the surface spectral function arising from the projection of the 3D-bulk spectral function on the (0-11) surface. Once more this component will not show any of the novel physics associated with the semi-infinite character of the system.
@@ II.4 QPI calculation
The QPI pattern corresponds to the Fourier transform of the local density of states at a specific energy. We compute the QPI pattern resulting from the physical scenario of an impurity localized at the (0-11) surface. This can then directly be compared to STM experiments [13, 47], taking into account all the surface effects. To compute the QPI for such a point-like impurity, the surface Green function $G_{s}(E,\mathbf{k}_{\parallel})$ has first to be computed using Eq. (19). Secondly, one has to compute the T-matrix associated to the point-like impurity [33], whose poles correspond to impurity states
$T_{s}(E)=\left(\mathds{1}-\mathds{U}_{s}\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}G_{s}(E,\mathbf{k}_{\parallel})\right)^{-1}\mathds{U}_{s}$ (24)
@@
$\mathds{U}_{s}=U_{0}\begin{pmatrix}\mathds{1}_{8}&0\\ 0&-\mathds{1}_{8}\\ \end{pmatrix}$ (25)
with $U_{0}$ the point-like impurity strength. The physical observables, such as the local density of states (LDOS) that can be measured near an impurity, can be expressed directly in terms of this T-matrix, if we assumes the dilute-limit approximation, such that the impurities are well separated from each other. The Fourier transform of the change in LDOS induced by the impurity, $\delta\rho(E,\mathbf{q}_{\parallel})$, can then be written as
$\delta\rho(E,\mathbf{q}_{\parallel})=-\dfrac{1}{2\pi i}\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}\mathrm{Tr}_{\mathrm{el}}\left[\,\widetilde{g}(E,\mathbf{k}_{\parallel},\mathbf{q}_{\parallel})\right]$
@@
-- ${G}_{s}^{\ast}(E,\mathbf{k}_{\parallel}+\mathbf{q}_{\parallel})T_{s}^{\ast}(E){G}_{s}^{\ast}(E,\mathbf{k}_{\parallel})$
@@
It is worth noting that the T-matrix method described here is exact, since it takes into into account the summation to all orders in the impurity strength. Simpler alternative to the T-matrix are the Born approximation [33], where only the first order scattering is considered, as well as the joint density of states (JDOS) technique, for which the Fourier transform of the LDOS is predicted to be described by an auto-correlation of the surface spectral function [48]:
$J(E,\mathbf{q}_{\parallel})=\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}{A}_{s}(E,\mathbf{k}_{\parallel}){A}_{s}(E,\mathbf{k}_{\parallel}+\mathbf{q}_{\parallel}).$ (28)
However, in general, the QPI patterns based on the T-matrix and JDOS calculations do not coincide. This is because the phase factors in the Green’s function matrix are not taken into account in the JDOS[49]. The QPI pattern computed with Eq. (II.4) is taking these interference effects into account.
@@ III Results for normal UTe2
We start our study by considering UTe2 in its normal state, thus setting the superconducting parameter to zero, $\Delta_{0}=0$. The parameters used for the tight-binding model are given in Appendix A.1. We fix the value $\eta=0.1$ meV for the quasiparticle damping, consistent with the experimental observations in Ref. 1. The value of the impurity potential simulating the hard edges is taken to be $V_{0}=1000$ eV, however any value at least one order of magnitude above the bandwidth would suffice [32]. The value for the local impurity potential used in the QPI calculation is $U_{0}=0.2$ eV. The sampling $(N_{k_{\#}}\times N_{k_{x}})$ of the $\mathbf{k}$-space along the (0-11) plane is $(200\times 200)$, and we use for the integration along the direction perpendicular to the plane (0-11) the value $N_{k_{\perp}}=5000$, which ensures a convergence of the QPI numerical results.
@@ IV Results for superconducting UTe2
In this Section we study UTe2 in its superconducting state for both non-chiral and chiral superconducting pairings. We first present the results for the density of states, and then we describe the surface spectral function and the QPI patterns. The value of the superconducting parameter is set to $\Delta_{0}=0.3$ meV, and the quasiparticle damping to $\eta=0.1$ meV, in agreement with the value of the thermal damping typically observed in experiments[1]. The tight-binding parameters, given in Appendix A.1, and the sampling of the $\mathbf{k}$-momentum are unchanged compared to the normal case.
@@ IV.1 Density of states
$\rho_{0}(E)$ $=$ $-\dfrac{1}{\pi}\int_{BZ}\text{Im}\{\text{Tr}\big{[}G_{\mathrm{BdG}}(E,\mathbf{k})\big{]}\}\mathrm{d}^{3}\mathbf{k}$ (29)
$\rho_{s}(E)$ $=$ $-\dfrac{1}{\pi}\int_{BZ_{\parallel}}\text{Im}\{\text{Tr}\big{[}G_{s}(E,\mathbf{k}_{\parallel})\big{]}\}\mathrm{d}^{2}\mathbf{k}_{\parallel}$ (30)
where $G_{\mathrm{BdG}}(E,\mathbf{k})$ and $G_{s}(E,\mathbf{k}_{\parallel})$ are the Bogoliubov-de-Gennes and respectively the surface Green’s functions given by Eqs. (14) and (19). The integral in Eq. 29 is performed over the entire 3D BZ. The results are presented in Figs. 4 and 5 for all the pairing symmetries considered in this study. The overall trend is that non-chiral pairings exhibit a U-shape bulk density of states while the chiral ones show a V-shape bulk density of states. We note also that the $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ pairing symmetries exhibit a maximal value at zero energy in the (0-11) surface density of states, indicating the formation of surface states in the vicinity of zero energy, while the other pairings have a minimal value at zero energy and two maxima at subgap energies of the order of $E\approx\pm 0.2$ meV, indicating the formation of surface states close to this energy value.
Figure 4: Density of states for non-chiral pairings. The black curve represents the bulk density of states, while the red curve corresponds to the surface density of states in the presence of a (0-11) impurity plane which mimics a surface in a 3D sample.
Figure 5: Same as in Fig. 4 for chiral pairings.
@@ IV.2 Bulk spectral function
[…]
@@
Figure 6: Bulk spectral function in the (0-11)-plane for non-chiral pairings at $E=0$. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
Figure 7: Same as in Fig. 6 for chiral pairings.

### v2

@@ II.3 Surface Green function calculation
The $T$-matrix method is an exact analytical method to compute Green’s function for a $\delta$-localized impurity in a non-interacting infinite system [balatsky2006impurity]. This method can be extended to compute the boundary states of a $d$-dimensional bulk Hamiltonian by considering a (dd-1)-dimensional localized impurity, which effectively splits the system in half [kaladzhyan2019obtaining, pinon2020surface] for a very large impurity potential. For example, for a 3D system, the Green’s function computed near a plane-like impurity will converge to the surface Green function in the limit of a large impurity potential.
@@
$G_{\mathrm{BdG}}(E,\mathbf{k})=\Big((E+i\eta)\mathbb{1}_{16}-\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})\Big)^{-1}~,$ (14)
with $\eta$ the quasiparticle damping. $\mathcal{H}_{\mathrm{BdG}}(\mathbf{k})$ is a $16\times 16$ matrix in the Hilbert space $\mathcal{E}=\mathcal{E}_{orb}\otimes\mathcal{E}_{spin}\otimes\mathcal{E}_{el-h}$ with $\mathcal{E}_{orb}$ the orbital subspace $\mathcal{E}_{spin}$ the spin subspace and $\mathcal{E}_{el-h}$ the electron-hole subspace (see Appendix A.2). $\mathbb{1}_{n}$ is an $n\times n$ identity matrix. To compute surface states in the cleave plane (0-11), we consider an impurity plane such that translation invariance is maintained in direction $\mathbf{k}_{\parallel}$ parallel to the plane and broken in orthogonal direction $\mathbf{k}_{\perp}$. By decomposing $\mathbf{r}$ into vectors perpendicular and parallel to the impurity plane, $\mathbf{r}=(\mathbf{r}_{\parallel},\mathbf{r}_{\perp})$, the impurity potential is given by $V(\mathbf{r})=\mathds{V}\delta(\mathbf{r}_{\perp})$, with
$\mathds{V}=V_{0}\begin{pmatrix}\mathds{1}_{8}&0\\ 0&-\mathds{1}_{8}\\ \end{pmatrix}~$, (15)
@@
$T(E,\mathbf{k}_{\parallel})=\left(\mathds{1}_{16}-\mathds{V}\int_{BZ_{\perp}}\dfrac{\mathrm{d}\mathbf{k}_{\perp}}{L_{BZ_{\perp}}}G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp},\mathbf{k}_{\parallel})\right)^{-1}\mathds{V}~,$
where $L_{BZ_{\perp}}$ is the length of the Brillouin zone along the axis perpendicular to the (0-11) plane, and where we have introduced $\mathbf{k}=(\mathbf{k}_{\parallel},\mathbf{k}_{\perp})$. The technical details on the $\mathbf{k}_{\perp}$ integration are given in Appendix A.3. From the knowledge of the T-matrix, the Green function in the presence of the impurity plane can be exactly computed since one has
@@
$+G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp 1},\mathbf{k}_{\parallel})T(E,\mathbf{k}_{\parallel})G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp 2},\mathbf{k}_{\parallel})~.$ (17)
@@
$\times G(E,\mathbf{k}_{\perp 1},\mathbf{k}_{\perp 2},\mathbf{k}_{\parallel})~,$
with $d_{\perp}=bc/\sqrt{b^{2}+c^{2}}$ is the distance between the (0-11) impurity plane and its nearest lattice plane. Using Eq. II.3, we can decompose $G_{s}(E,\mathbf{k}_{\parallel})$ into a sum of two terms
$G_{s}(E,\mathbf{k}_{\parallel})=G_{b}(E,\mathbf{k}_{\parallel},0)+G_{i}(E,\mathbf{k}_{\parallel})~,$ (19)
@@
$G_{b}(E,\mathbf{k}_{\parallel},z_{\perp})$ $=$ $\int_{BZ_{\perp}}\dfrac{\mathrm{d}\mathbf{k}_{\perp}}{L_{BZ_{\perp}}}e^{iz_{\perp}\mathbf{k}_{\perp}\cdot\mathbf{e}_{\perp}}G_{\mathrm{BdG}}(E,\mathbf{k}_{\perp},\mathbf{k}_{\parallel})~,$
$G_{i}(E,\mathbf{k}_{\parallel})$ $=$ $G_{b}(E,\mathbf{k}_{\parallel},d_{\perp})T(E,\mathbf{k}_{\parallel})G_{b}(E,\mathbf{k}_{\parallel},-d_{\perp})~.$
We have chosen an integration region with the same area as the first BZ, but of a rectangular shape, following Ref. [pinon2020surface]. Due to the $k$-space periodicity, this is equivalent to integrating over the first BZ and much more straightforward. We note that the bulk contribution corresponds solely to a simple projection of the 3D-bulk physics on the (0-11) surface, without taking into account the existence of a surface and the semi-infinite nature of the system. The surface physics is hidden in the impurity term $G_{i}(E,\mathbf{k}_{\parallel})$ which when the impurity potential is taken to infinity will describe the formation of the surface states.
@@
$A_{s}(E,\mathbf{k}_{\parallel})=-\frac{1}{\pi}\mathrm{Im}\{\mathrm{Tr}_{\mathrm{el}}[G_{s}(E,\mathbf{k}_{\parallel})]\}~,$ (22)
where the trace runs over the electron bands. We can also here distinguish two components, and in what follows we sometimes calculate separately the “bulk” spectral function
$A_{b}(E,\mathbf{k}_{\parallel})=-\frac{1}{\pi}\mathrm{Im}\{\mathrm{Tr}_{\mathrm{el}}[G_{b}(E,\mathbf{k}_{\parallel},0)]\}~,$ (23)
which is just the component of the surface spectral function arising from the projection of the 3D-bulk spectral function on the (0-11) surface. Once more, this component will not show any of the novel physics associated with the semi-infinite character of the system.
@@ II.4 QPI calculation
The QPI pattern corresponds to the Fourier transform of the local density of states at a specific energy. We compute the QPI pattern resulting from the physical scenario of an impurity located on the (0-11) surface. This can then be directly compared with STM experiments [gu2023detection, gu2025pair], taking into account all the surface effects. To compute the QPI for such a point-like impurity, the surface Green function $G_{s}(E,\mathbf{k}_{\parallel})$ has first to be computed using Eq. (19). Second, one has to compute the T-matrix associated with the point-like impurity [balatsky2006impurity], whose poles correspond to impurity states
$T_{s}(E)=\left(\mathds{1}-\mathds{U}_{s}\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}G_{s}(E,\mathbf{k}_{\parallel})\right)^{-1}\mathds{U}_{s}~,$ (24)
@@
$\mathds{U}_{s}=U_{0}\begin{pmatrix}\mathds{1}_{8}&0\\ 0&-\mathds{1}_{8}\\ \end{pmatrix}~$, (25)
with $U_{0}$ the strength of the point-like impurity. The scatterer considered in our analysis is non-magnetic and has equal components on all four orbitals. We checked that QPI features are qualitatively unchanged for different forms of the potential. The physical observables, such as the local density of states (LDOS) that can be measured near an impurity, can be expressed directly in terms of this T-matrix, if we assume the dilute-limit approximation, such that the impurities are well separated from each other. The Fourier transform of the change in LDOS induced by the impurity, $\delta\rho(E,\mathbf{q}_{\parallel})$, can then be written as
$\delta\rho(E,\mathbf{q}_{\parallel})=-\dfrac{1}{2\pi i}\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}\mathrm{Tr}_{\mathrm{el}}\left[\,\widetilde{g}(E,\mathbf{k}_{\parallel},\mathbf{q}_{\parallel})\right]~,$
@@
- ${G}_{s}^{\ast}(E,\mathbf{k}_{\parallel}+\mathbf{q}_{\parallel})T_{s}^{\ast}(E){G}_{s}^{\ast}(E,\mathbf{k}_{\parallel})~.$
@@
It is worth noting that the T-matrix method described here is exact, since it takes into account the summation to all orders in the impurity strength. Simpler alternative to the T-matrix is the Born approximation [balatsky2006impurity], where only the first-order scattering is considered, as well as the joint density of states (JDOS) technique, for which the Fourier transform of the LDOS is predicted to be described by an auto-correlation of the surface spectral function [roushan2009topological]:
$J(E,\mathbf{q}_{\parallel})=\int_{BZ_{\parallel}}\dfrac{\mathrm{d}^{2}\mathbf{k}_{\parallel}}{S_{BZ_{\parallel}}}{A}_{s}(E,\mathbf{k}_{\parallel}){A}_{s}(E,\mathbf{k}_{\parallel}+\mathbf{q}_{\parallel})~.$ (28)
However, in general, the QPI patterns based on the T-matrix and JDOS calculations do not coincide. This is because the phase factors in the Green’s function matrix are not taken into account in the JDOS[fang2013theory]. The QPI pattern computed with Eq. (II.4) is taking these interference effects into account.
@@ III Results for normal UTe2
We start our study by considering UTe2 in its normal state, thus setting the superconducting parameter to zero: $\Delta_{0}=0$. The parameters used for the tight-binding model are given in Appendix A.1. We fix the value $\eta=0.1$ meV for the quasiparticle damping, consistent with the experimental observations in Ref. experimentalpaper. The value of the impurity potential used to simulate hard edges is taken to be $V_{0}=1000$ eV, however, any value at least one order of magnitude above the bandwidth would suffice [pinon2020surface]. The value of the local impurity potential used in the QPI calculation is $U_{0}=0.2$ eV. The sampling $(N_{k_{\#}}\times N_{k_{x}})$ of the $\mathbf{k}$-space along the (0-11) plane is $(200\times 200)$, and we use the value $N_{k_{\perp}}=5000$ for integration along the direction perpendicular to the plane (0-11) to ensure convergence of the QPI numerical results.
@@ IV Results for superconducting UTe2
In this Section we study UTe2 in its superconducting state for both non-chiral and chiral superconducting pairings. We first present the results for the density of states and then describe the surface spectral function and the QPI patterns. The value of the superconducting parameter is set to $\Delta_{0}=0.3$ meV, and the quasiparticle damping to $\eta=0.1$ meV, in agreement with the value of the thermal damping typically observed in experiments[experimentalpaper]. The tight-binding parameters, given in Appendix A.1, and the sampling of the $\mathbf{k}$-momentum are unchanged compared with the normal case.
@@ IV.1 Density of states
$\rho_{0}(E)$ $=$ $-\dfrac{1}{\pi}\int_{BZ}\text{Im}\{\text{Tr}\big[G_{\mathrm{BdG}}(E,\mathbf{k})\big]\}\mathrm{d}^{3}\mathbf{k}~,$ (29)
$\rho_{s}(E)$ $=$ $-\dfrac{1}{\pi}\int_{BZ_{\parallel}}\text{Im}\{\text{Tr}\big[G_{s}(E,\mathbf{k}_{\parallel})\big]\}\mathrm{d}^{2}\mathbf{k}_{\parallel}~,$ (30)
where $G_{\mathrm{BdG}}(E,\mathbf{k})$ and $G_{s}(E,\mathbf{k}_{\parallel})$ are the Bogoliubov-de-Gennes and respectively the surface Green’s functions given by Eqs. (14) and (19). The integral in Eq. 29 is performed over the entire 3D BZ. The results are presented in Figs. 4 and 5 for all the pairing symmetries considered in this study. All chiral and non-chiral pairings have a U-shaped bulk density of states at low energies, consistent with point nodes [ishihara2023chiral]. Some pairings acquire a V-shape at larger energies when other bands start contributing.
We note also that the $B_{3u}$, $A_{u}+iB_{1u}$ and $A_{u}+iB_{2u}$ pairing symmetries exhibit a maximal value at zero energy in the (0-11) surface density of states, indicating the formation of surface states in the vicinity of zero energy, while the other pairings have a minimal value at zero energy and two maxima at subgap energies of the order of $E\approx\pm 0.2$ meV, indicating the formation of surface states close to this energy value. The DOS profile can be affected by additional allowed symmetry d-vector contributions, as discussed in Appendix A.4. In particular, the zero-energy peak observed for the $B_{3u}$ order parameter can split upon adding additional d-vector terms.
Figure 4: Density of states for non-chiral pairings. The black curve represents the bulk density of states, while the red curve corresponds to the surface density of states in the presence of a (0-11) impurity plane which mimics a surface in a 3D sample. The horizontal gray line corresponds to the DOS in the normal state. At low energies, the bulk density of states for all non-chiral pairings exhibits a characteristic U-shaped profile, consistent with either gapped or nodal superconductivity in three dimensions. The $B_{3u}$ pairing is showing a pronounced zero-energy peak in the surface DOS in contrast with the other pairings still exhibiting a U-shaped profile. We have $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
Figure 5: Density of states for chiral pairings. The black curve represents the bulk density of states, while the red curve corresponds to the surface density of states in the presence of a (0-11) impurity plane which mimics a surface in a 3D sample. The horizontal gray line corresponds to the DOS in the normal state. At low energies, the bulk density of states for all chiral pairings exhibits a characteristic U-shaped profile, consistent with either gapped or nodal superconductivity in three dimensions. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.
@@ IV.2 Bulk spectral function
The bulk contribution to the spectral function $A_{b}(E,k_{\#},k_{x})$, defined by Eq. 23, essentially represents the direct projection of the imaginary part of the bulk Green function onto the considered surface, i.e. (0-11)- plane. Figures 6 and   7 show this quantity at zero-energy, for both non-chiral and chiral pairings. Note that the overall amplitude for $A_{b}(E,k_{\#},k_{x})$ is reduced when UTe2 is in its superconducting state compared with UTe2 in its normal state. Apart from this amplitude reduction, we observe that the bulk spectral function profile as a function $k_{\#}$ and $k_{x}$ remains generally similar to the normal state one (see Fig. 3). However, as discussed in detail in Ref.  experimentalpaper, for some superconducting order parameters, in particular $B_{2u}$ and $B_{3u}$ among the non-chiral order parameters, one expects the presence of extra features, in particular of nodal points. Here, some of these nodes are masked by significant quasiparticle damping, chosen to match the realistic experimental conditions. However, they can become visible for significantly smaller damping, as illustrated in Fig. 21 in Appendix B.5.
@@
Figure 6: Bulk spectral function in the (0-11)-plane for non-chiral pairings at $E=0$. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV. The pairings exhibit similar spectral functions, with differences arising from their distinct nodal structures.
Figure 7: Bulk spectral function in the (0-11)-plane for chiral pairings at $E=0$. We take $\Delta_{0}=0.3$ meV and $\eta=0.1$ meV.